# Controlling Nonequilibrium Bose-Einstein Condensation with Engineered Environments

### v3

@@ VI Values of the simulation parameters and further examples
For the resonator frequencies $\omega_{j}$, the exact value is not crucial but what is important is that (i) they differ by an amount of the order of the tunnelling coupling $J$, such that the eigenstates are delocalized over several resonators; (ii) accidentally identical gaps in the spectrum are avoided. Similar tigh-binding arrays have been realized (also with additional nonlinearities to implement photon-photon interactions, which are not needed here), for instance, in Refs. [41, 30, 51, 31, 32]. We choose arbitrarily $\omega_{j}=\omega_{R}+J(m/10+\varepsilon_{m})$, relative to a common frequency $\omega_{R}$, where $\varepsilon_{m}\in[0,1)$ is a disorder parameter picked from one realization of uniformly-distributed random numbers, specifically
@@
rounded to two significant digits. Although our simulations only depend on the ratio $\omega_{j}/J$, realistic values of $\omega_{R}$ and $J$ may be chosen, for instance, as $\omega_{R}/2\pi=5$ GHz and $J/2\pi=30$ MHz [29]. The resulting spectrum for the five-site chain is shown in Fig. 2(a). All reservoir spins decay strongly with rate $\gamma_{\ell}=0.2J$. For the choice of $J/2\pi=30$ MHz, this gives $\gamma_{\ell}/2\pi=6$ MHz, compatible with values implemented for reservoir transmons, e.g., in Ref. [32]. The parameter $\delta_{\ell}$ is fixed by (i) enforcing that the $\ell$th-spin level splitting $\mathcal{E}_{\ell}=|\delta_{\ell}|\sqrt{1+r_{\ell}^{2}}$ resonates with a chosen system level distance $E_{ij}$ of the bosonic chain, and (ii) imposing the sign $\mathrm{sgn}(\delta_{\ell})$ found from the solution $\bm{z}$ of the inequalities of Eq. (5) for a given target condensation pattern. The values of the parameter $r_{\ell}$ are then chosen such that the Rabi frequency $\delta_{\ell}r_{\ell}$ of the microwave drives on the qubits meets experimental constraints—in particular, such that it remains around an order of magnitude smaller than the transmon nonlinearities, which have typical values of $\sim 300$ MHz, to prevent unwanted excitations outside of the lowest two-level subspace. For the five-site chain with connectivity of Fig. 2(a), up to the sign of $\delta_{\ell}$, which depends on the choice of condensation patterns, the values used are then (up to three significant digits)
@@ VII Further experimental considerations
Temperature effects and photon emission. At the dilution-refrigerator temperatures $\sim 10$ mK at which superconducting circuits operate, the absorption of thermal photons is strongly suppressed and would only contribute a small addition of particles to the Bose selection. Considering superconducting microwave resonators with routinely-achieved $Q$-factors of $Q\sim 10^{5}-10^{6}$, the photon decay rate $\kappa/2\pi\sim 1-10$ kHz by spontaneous emission are one to two orders of magnitude smaller than the dissipation engineered through the reservoir spins. The same applies to weak losses potentially originating from Purcell-like decay due to the weak hybridization of the resonators with the strongly damped reservoir spins, which can be estimated as $(g_{\ell}/\Delta_{\ell})^{2}\gamma_{\ell}$ [29] (with $g$ and $\Delta$ defined in Sec. VI and $\gamma_{\ell}$ the spin decay rate). The net particle loss will thus be weak within the timescales necessary for the system to reach its engineered Bose-selected steady state. Also, as discussed in the main text and in detail in Sec. VIII, Bose selection is driven by terms in the mean-field Eq. (2), which are quadratic in the large occupations $n_{i}$, whereas losses by photon emission depend linearly on $n_{i}$. Therefore, ‘pre-relaxation’ to the Bose selected state occurs much faster than the photon leakage. Finally, the system can be fully stabilized also against this weak photon loss, such that losses are sufficiently compensated and do not alter the Bose selection pattern also at long times, by using additional qubit reservoirs as particle pumps. This is detailed below, in Sec. VIII.
Interactions. The perfectly harmonic spectrum of the microwave resonators provides a system hosting non-interacting photons, as desired to study Bose selection driven purely by quantum statistics. Spurious photon-photon interactions may in principle arise due to hybridization with the ancillary reservoir spins, inducing effective nonlinearities in the resonator array. However, since the spins operate in the dispersive regime, the latter (as well as resonant photon exchange yielding particle loss) can be neglected. This is, in a way, substantiated by the fact that tracing out the spins in favour of effective quantum-jump rates $R_{ij}$ among the eigenstates of the resonators system yields a very good description, as shown quantitatively in Fig. 2(d). Spin-spin interaction mediated by the resonators, potentially influencing the engineered-dissipation, may arise in case the qubits happen to have the same transition frequency [53], but this will in general require fine-tuning, and thus also this effect is not expected to play a significant role in the proposed implementation.
@@ VIII Impact of pump and loss
$R_{n+1,n}$ $=\left(\frac{g}{2}\right)^{2}|\!\bra{n+1}\hat{a}^{\dagger}\ket{n}\!|^{2}S(0)$ (S85)

### v4

@@ VI Values of the simulation parameters and further examples
For the resonator frequencies $\omega_{\ell}$, the exact value is not crucial but what is important is that (i) they differ by an amount of the order of the tunnelling coupling $J$, such that the eigenstates are delocalized over several resonators; (ii) accidentally identical gaps in the spectrum are avoided. Similar tight-binding arrays have been realized (also with additional nonlinearities to implement photon-photon interactions, which are not needed here), for instance, in Refs. [42, 30, 54, 31, 32]. We choose arbitrarily $\omega_{\ell}=\omega_{R}+J(\ell/10+\varepsilon_{\ell})$, relative to a common frequency $\omega_{R}$, where $\varepsilon_{\ell}\in[0,1)$ is a disorder parameter picked from one realization of uniformly-distributed random numbers, specifically
@@
rounded to two significant digits. Although our simulations only depend on the ratio $\omega_{\ell}/J$, realistic values of $\omega_{R}$ and $J$ may be chosen, for instance, as $\omega_{R}/2\pi=5$ GHz and $J/2\pi=30$ MHz [29]. The resulting spectrum for the five-site chain is shown in Fig. 2(a). All reservoir spins decay strongly with rate $\gamma_{\ell}=0.2J$. For the choice of $J/2\pi=30$ MHz, this gives $\gamma_{\ell}/2\pi=6$ MHz, compatible with values implemented for reservoir transmons, e.g., in Ref. [32]. The parameter $\delta_{\ell}$ is fixed by (i) enforcing that the $\ell$th-spin level splitting $\mathcal{E}_{\ell}=|\delta_{\ell}|\sqrt{1+r_{\ell}^{2}}$ resonates with a chosen system level distance $E_{ij}$ of the bosonic chain, and (ii) imposing the sign $\mathrm{sgn}(\delta_{\ell})$ found from the solution $\bm{z}$ of the inequalities of Eq. (7) for a given target condensation pattern. The values of the parameter $r_{\ell}$ are then chosen such that the Rabi frequency $\delta_{\ell}r_{\ell}$ of the microwave drives on the qubits meets experimental constraints—in particular, such that it remains around an order of magnitude smaller than the transmon nonlinearities, which have typical values of $\sim 300$ MHz, to prevent unwanted excitations outside of the lowest two-level subspace. For the five-site chain with connectivity of Fig. 2(a), up to the sign of $\delta_{\ell}$, which depends on the choice of condensation patterns, the values used are then (up to three significant digits)
@@ VII Further experimental considerations
Temperature effects and photon emission. At the dilution-refrigerator temperatures $\sim 10$ mK at which superconducting circuits operate, the absorption of thermal photons is strongly suppressed and would only contribute a small addition of particles to the Bose selection. Considering superconducting microwave resonators with routinely-achieved $Q$-factors of $Q\sim 10^{5}-10^{6}$, the photon decay rate $\kappa/2\pi\sim 1-10$ kHz by spontaneous emission are one to two orders of magnitude smaller than the dissipation engineered through the reservoir spins. The same applies to weak losses potentially originating from Purcell-like decay due to the weak hybridization of the resonators with the strongly damped reservoir spins, which can be estimated as $(g_{\ell}/\Delta_{\ell})^{2}\gamma_{\ell}$ [29] (with $g$ and $\Delta$ defined in Sec. VI and $\gamma_{\ell}$ the spin decay rate). The net particle loss will thus be weak within the timescales necessary for the system to reach its engineered Bose-selected steady state. Also, as discussed in the main text and in detail in Sec. VIII, Bose selection is driven by terms in the mean-field Eq. (S12), which are quadratic in the large occupations $n_{i}$, whereas losses by photon emission depend linearly on $n_{i}$. Therefore, ‘pre-relaxation’ to the Bose selected state occurs much faster than the photon leakage. Finally, the system can be fully stabilized also against this weak photon loss, such that losses are sufficiently compensated and do not alter the Bose selection pattern also at long times, by using additional qubit reservoirs as particle pumps. This is detailed below, in Sec. VIII.
Interactions. The perfectly harmonic spectrum of the microwave resonators provides a system hosting non-interacting photons, as desired to study Bose selection driven purely by quantum statistics. Spurious photon-photon interactions may in principle arise due to hybridization with the ancillary reservoir spins, inducing effective nonlinearities in the resonator array. However, since the spins operate in the dispersive regime, the latter (as well as resonant photon exchange yielding particle loss) can be neglected. This is, in a way, substantiated by the fact that tracing out the spins in favour of effective quantum-jump rates $R_{ij}$ among the eigenstates of the resonators system yields a very good description, as shown quantitatively in Fig. 2(d). Spin-spin interaction mediated by the resonators, potentially influencing the engineered-dissipation, may arise in case the qubits happen to have the same transition frequency [Majer2007], but this will in general require fine-tuning, and thus also this effect is not expected to play a significant role in the proposed implementation.
@@ VIII Impact of pump and loss
$R_{n+1,n}$ $=\left(\frac{g}{2}\right)^{2}|\!\bra{n+1}\hat{a}^{\dagger}\ket{n}\!|^{2}S(0)\ ,$ (S86)